OpenStudy (anonymous):

How do you rewrite the expression and simplify (1/3 -2b)27 using distributive property

OpenStudy (anonymous):

distributive property is \[ a\times(b+c) = a\times b+ a\times c\]So the given expression becomes\[(\frac{1}{3}- 2b)27= \frac{1}{3}\times27 - 2b\times 27\]simplify to get the final answer.